OpenStudy (anonymous):

prove y=(ln(x) + c)/x is a soln. of: x^2y' + xy = 1

OpenStudy (anonymous):

Could you post again your question? please be more clear.

OpenStudy (anonymous):

\[y=(\ln(x) + c)/x\] solution of: \[x^2y' + xy = 1\]

OpenStudy (anonymous):

prove it; \[y=(\ln(x)+c)/x\] \[xy=\ln(x)+c\] implicit derivative \[xy'+y=1/x\] multiply for x to both sides \[x^2y′+xy=1\]

OpenStudy (anonymous):

could you became my fan!

OpenStudy (apples):

\[x^2y' + xy = 1\]\[x^2 \cdot -\frac{c + \log{x} - 1}{x^2} + x\frac{\log{x} + c}{x} = 1\]\[\log{x} + c - (c+\log{x}-1) = 1\]\[1=1\] Thus \[\frac{\log{x}+c}{x}\] is a general solution of that differential equation.